OpenStudy (anonymous):

Solve 2 log x = log 64. x = 8 x = ±8 x = 32 x = 128

OpenStudy (campbell_st):

rewrite the equation using index laws \[\log(x^2) = \log(64)\] so now you have what appears to be the same base... then you can say \[x^2 = 64\] solve for x and alternative is \[2\log(x) = \log(8^2)\] and then using log laws \[2\log(x) = 2 \log(8)\] you need to think about the solution can you find the log of a negative

OpenStudy (anonymous):

idk how to solve the problem

OpenStudy (campbell_st):

well look at log(x) = log(8) what do you think is the value of x....?

OpenStudy (anonymous):

8?

OpenStudy (campbell_st):

yes... that's the answer

OpenStudy (anonymous):

so simple. thank you